# Reconfigurable routing in data center networks

### v1

@@ 5 Proofs
Proof.
Note that RRP as in the statement of the theorem is in NP on account of the numerical restrictions imposed. As in the proof of Theorem 9, we reduce from the problem RXC3. Let $(\mathcal{X},\mathcal{C})$ be an instance of RXC3 of size $n$. W.l.o.g., we may assume that $n$ is even. Let $m=\lceil\log_{2}(3n)\rceil$ (so, $3n\leq 2^{m}$). Our static network $H$ is the hypercube $Q_{8m}$ (so, $|Q_{8m}|=2^{8m}=O(n^{8})$) and there is exactly $1$ switch link from every node to our switch $s$. We begin by defining a weighted digraph $D^{\prime}=(V,E^{\prime})$ where $V$ is the node set of our hypercube $Q_{8m}$. The digraph $D^{\prime}$ is to provide a description of our workload matrix. In order to define the directed edges of $D^{\prime}$, we define some specific subsets of nodes of $V$. We explain the notation that we use as we proceed. By ‘distance’ we mean the distance (that is, shortest path length) in the hypercube $Q_{8m}$ and by $V_{i}$ (resp. $V_{\leq i}$) we mean the subset of nodes of $V$ that are at distance exactly $i$ (resp. at most $i$) from the root node $r=(1,1,\ldots,1)=1^{8m}$ (in general, we also denote nodes of $V$ as bit-strings of length $8m$).
  • 1.
@@
Proof.
Define a perfect matching $M$ on $W$ as follows. First, match every node in $V_{4m}$ with the node $v$ that differs from $u$ in every bit; that is, $v=\bar{u}$. Consider some node $z\bar{z}z\bar{z}z\bar{z}z\bar{z}$ of $X$. Its matched node is $\bar{z}z\bar{z}z\bar{z}z\bar{z}z$ which is either in $X$ or outside $X$. Suppose it is the latter. Suppose also that some other node $w\bar{w}w\bar{w}w\bar{w}w\bar{w}$ is in $X$ and its matched node $\bar{w}w\bar{w}w\bar{w}w\bar{w}w$ lies outside $X$. Note that the distance between the two matched nodes $\bar{z}z\bar{z}z\bar{z}z\bar{z}z$ and $\bar{w}w\bar{w}w\bar{w}w\bar{w}w$ is at least $8$. We amend our matching $M$ by removing all matched pairs of nodes of $X$ and choose an arbitrary matching on the remaining matched nodes (note that there is an even number of such matched nodes). Any pair of nodes in $M$ is such that there is no static link joining them and every node of $V_{4m}\setminus X$ is involved in $M$.
@@
In order to complete our instance $(H(S),D,\mu,\kappa)$, we define $\mu$ to be any fixed (rational) value in the interval $(0,1)$ and $\kappa=\kappa_{\beta}+\kappa_{\alpha}+\kappa_{1}$ where:
  • 1.
@@
This constitutes the configuration $N$. Consider some $\beta$-demand of our workload. For some node $c_{j}\in C^{\prime}$, all such workloads originating at the nodes $c_{j}^{\oplus 001}$, $c_{j}^{\oplus 010}$ and $c_{j}^{\oplus 100}$ can be served via a flow-path consisting of a solitary dynamic link at workload cost $\mu\beta$. For some node $c_{j}\in C\setminus C^{\prime}$, suppose that there is a demand $D[c_{j}^{\oplus 001},x_{1}^{i}]=\beta$. This demand exists because the element $x_{i}$ is the first element of clause $c_{j}$. As the element $x_{i}$ is the first element of clause $c_{j}$, there is a dynamic link from $c_{j}^{\oplus 001}$ to a neighbour of $x_{i}$ in $\{x_{i}^{\oplus 01},x_{i}^{\oplus 10}\}$, which w.l.o.g. we may assume to be $x_{i}^{\oplus 01}$; hence, the demand can be served via the flow-path $c_{j}^{\oplus 001},x_{i}^{\oplus 01},x_{i}$ at workload cost $(\mu+1)\beta$. Defining other flow-paths analogously means that we can find flow-paths corresponding to the directed edges of $E_{\beta}$ the total workload cost of which is $3n\mu\beta+6n(\mu+1)\beta=\kappa_{\beta}$. All $\alpha$-demands can be served via a flow-path consisting of a solitary dynamic link at total workload cost $|E_{\alpha}|\mu\alpha=\kappa_{\alpha}$. Consider some $1$-demand $D[r,x_{i}]=1$. Let the clause $c_{j}\in\mathcal{C}^{\prime}$ be such that $x_{i}$ is an element of $c_{j}$; hence, there is a dynamic link from some port node $p_{j^{\prime}}$ of $P$ to $c_{j}$ and a dynamic link from some neighbour of $c_{j}$ from $\{c_{j}^{\oplus 001},c^{\oplus 010},c^{\oplus 110}\}$ to $x_{i}$. Consequently, there is a static path from $r$ to $p_{j^{\prime}}$ of weight $m$, a dynamic link from $p_{j^{\prime}}$ to the node $c_{j}$, a static link from $c_{j}$ to the above neighbour in $\{c_{j}^{\oplus 001},c^{\oplus 010},c^{\oplus 110}\}$ and a dynamic link from this neighbour to $x_{i}$. Hence, the demand $D[r,x_{i}]=1$ can be served via a flow-path of workload cost $m+1+2\mu$ with all $1$-demands served via flow-paths at a total workload cost $3n(m+1+2\mu)=\kappa_{1}$. Consequently, our instance $(H(S),D,\mu,\kappa)$ is a yes-instance of RRP. Conversely, suppose that$(H(S),D,\mu,\kappa)$ is a yes-instance of RRP and that $N$ is a configuration and $F$ a set of flow-paths witnessing that the total workload cost is at most $\kappa$.
Claim 2.
Every directed edge $(u,v)$ of $E_{\alpha}$ is necessarily such that $(u,v)\in N$ and the total workload cost of flow-paths serving the $\alpha$-demands is exactly $\kappa_{1}$.
Proof.

### v2

@@ 3 Results
Proof
Note that RRP as in the statement of the theorem is in NP on account of the numerical restrictions imposed. As in the proof of Theorem 3.2, we reduce from the problem RXC3. Let $(\mathcal{X},\mathcal{C})$ be an instance of RXC3 of size $n$. W.l.o.g., we may assume that $n$ is even. Let $m=\lceil\log_{2}(3n)\rceil$ (so, $3n\leq 2^{m}$). Our static network $H$ is the hypercube $Q_{8m}$ (so, $|Q_{8m}|=2^{8m}=O(n^{8})$) and there is exactly $1$ switch link from every node to our switch $s$.
We begin by defining a weighted digraph $D^{\prime}=(V,E^{\prime})$ where $V$ is the node set of our hypercube $Q_{8m}$. The digraph $D^{\prime}$ is to provide a description of our workload matrix. In order to define the directed edges of $D^{\prime}$, we define some specific subsets of nodes of $V$. We explain the notation that we use as we proceed. By ‘distance’ we mean the distance (that is, shortest path length) in the hypercube $Q_{8m}$ and by $V_{i}$ (resp. $V_{\leq i}$) we mean the subset of nodes of $V$ that are at distance exactly $i$ (resp. at most $i$) from the root node $r=(1,1,\ldots,1)=1^{8m}$ (in general, we also denote nodes of $V$ as bit-strings of length $8m$).
@@
Proof
Define a perfect matching $M$ on $W$ as follows. First, match every node in $V_{4m}$ with the node $v$ that differs from $u$ in every bit; that is, $v=\bar{u}$.
Consider some node $z\bar{z}z\bar{z}z\bar{z}z\bar{z}$ of $X$. Its matched node is $\bar{z}z\bar{z}z\bar{z}z\bar{z}z$ which is either in $X$ or outside $X$. Suppose it is the latter. Suppose also that some other node $w\bar{w}w\bar{w}w\bar{w}w\bar{w}$ is in $X$ and its matched node $\bar{w}w\bar{w}w\bar{w}w\bar{w}w$ lies outside $X$. Note that the distance between the two matched nodes $\bar{z}z\bar{z}z\bar{z}z\bar{z}z$ and $\bar{w}w\bar{w}w\bar{w}w\bar{w}w$ is at least $8$. We amend our matching $M$ by removing all matched pairs of nodes of $X$ and choose an arbitrary matching on the remaining matched nodes (note that there is an even number of such matched nodes). Any pair of nodes in $M$ is such that there is no static link joining them and every node of $V_{4m}\setminus X$ is involved in $M$.
@@
In order to complete our instance $(H,\mu,D,\kappa)$, we define $\mu$ to be any fixed (rational) value in the interval $(0,1)$ and $\kappa=\kappa_{\beta}+\kappa_{\alpha}+\kappa_{1}$ where:
@@
This constitutes the configuration $N$.
Consider some $\beta$-demand of our workload. For some node $c_{j}\in C^{\prime}$, all such workloads originating at the nodes $c_{j}^{\oplus 001}$, $c_{j}^{\oplus 010}$ and $c_{j}^{\oplus 100}$ can be served via a flow-path consisting of a solitary dynamic link at workload cost $\mu\beta$. For some node $c_{j}\in C\setminus C^{\prime}$, suppose that there is a demand $D[c_{j}^{\oplus 001},x_{1}^{i}]=\beta$. This demand exists because the element $x_{i}$ is the first element of clause $c_{j}$. As the element $x_{i}$ is the first element of clause $c_{j}$, there is a dynamic link from $c_{j}^{\oplus 001}$ to a neighbour of $x_{i}$ in $\{x_{i}^{\oplus 01},x_{i}^{\oplus 10}\}$, which w.l.o.g. we may assume to be $x_{i}^{\oplus 01}$; hence, the demand can be served via the flow-path $c_{j}^{\oplus 001},x_{i}^{\oplus 01},x_{i}$ at workload cost $(\mu+1)\beta$. Defining other flow-paths analogously means that we can find flow-paths corresponding to the directed edges of $E_{\beta}$ the total workload cost of which is $3n\mu\beta+6n(\mu+1)\beta=\kappa_{\beta}$.
All $\alpha$-demands can be served via a flow-path consisting of a solitary dynamic link at total workload cost $|E_{\alpha}|\mu\alpha=\kappa_{\alpha}$.
Consider some $1$-demand $D[r,x_{i}]=1$. Let the clause $c_{j}\in\mathcal{C}^{\prime}$ be such that $x_{i}$ is an element of $c_{j}$; hence, there is a dynamic link from some port node $p_{j^{\prime}}$ of $P$ to $c_{j}$ and a dynamic link from some neighbour of $c_{j}$ from $\{c_{j}^{\oplus 001},c^{\oplus 010},c^{\oplus 110}\}$ to $x_{i}$. Consequently, there is a static path from $r$ to $p_{j^{\prime}}$ of weight $m$, a dynamic link from $p_{j^{\prime}}$ to the node $c_{j}$, a static link from $c_{j}$ to the above neighbour in $\{c_{j}^{\oplus 001},c^{\oplus 010},c^{\oplus 110}\}$ and a dynamic link from this neighbour to $x_{i}$. Hence, the demand $D[r,x_{i}]=1$ can be served via a flow-path of workload cost $m+1+2\mu$ with all $1$-demands served via flow-paths at a total workload cost $3n(m+1+2\mu)=\kappa_{1}$. Consequently, our instance $(H,\mu,D,\kappa)$ is a yes-instance of 1-switched RRP($\sigma=3$).
Conversely, suppose it is the case that $(H,\mu,D,\kappa)$ is a yes-instance of 1-switched RRP($\sigma=3$) and that $N$ is a configuration and $F$ a set of flow-paths witnessing that the total workload cost is at most $\kappa$.
Claim 2
Every directed edge $(u,v)$ of $E_{\alpha}$ is necessarily such that $(u,v)\in N$ and the total workload cost of flow-paths serving the $\alpha$-demands is exactly $\kappa_{\alpha}$.
Proof